OpenStudy (anonymous):

If the sum of the measures of the interior angles of a convex polygon is 2160o, how many sides does the polygon have?

OpenStudy (anonymous):

Sum of interior angles of a polygon of n sides is equal to (n.180-360) given 180n-360=2160 180n=2160+360 hence n=13 sides

OpenStudy (anonymous):

Actually, N=14 \[2160+360=2520\] \[2520/180=14\]